U3 Company is considering three long-term capital investment proposals. Each investment has a useful life of 5 years. Relevant data on each project are as follows.

Project Bono Project Edge Project Clayton
Capital investment $168,000 $183,750 $210,000
Annual net income:
Year  1 14,700 18,900 28,350
        2 14,700 17,850 24,150
        3 14,700 16,800 22,050
        4 14,700 12,600 13,650
        5 14,700 9,450 12,600
Total $73,500 $75,600 $100,800


Depreciation is computed by the straight-line method with no salvage value. The company’s cost of capital is 15%. (Assume that cash flows occur evenly throughout the year.)

Compute the cash payback period for each project. (Round answers to 2 decimal places, e.g. 10.50.)

Project Bono enter the cash payback period in years rounded to 2 decimal places years
Project Edge enter the cash payback period in years rounded to 2 decimal places years
Project Clayton enter the cash payback period in years rounded to 2 decimal places years

Compute the net present value for each project. (Round answers to 0 decimal places, e.g. 125. If the net present value is negative, use either a negative sign preceding the number eg -45 or parentheses eg (45). For calculation purposes, use 5 decimal places as displayed in the factor table provided.)

Project Bono Project Edge Project Clayton
Net present value $enter a dollar amount rounded to 0 decimal places $enter a dollar amount rounded to 0 decimal places $enter a dollar amount rounded to 0 decimal places

Compute the annual rate of return for each project. (Hint: Use average annual net income in your computation.) (Round answers to 2 decimal places, e.g. 10.50%.)

Project Bono Project Edge Project Clayton
Annual rate of return enter percentages rounded to 2 decimal places % enter percentages rounded to 2 decimal places % enter percentages rounded to 2 decimal places %

Rank the projects on each of the foregoing bases. Which project do you recommend?

Project Cash Payback Net
Present Value
Annual
Rate of Return
Bono select a ranking number                                                                      132 select a ranking number                                                                      231 select a ranking number                                                                      321
Edge select a ranking number                                                                      132 select a ranking number                                                                      123 select a ranking number                                                                      123
Clayton select a ranking number                                                                      132 select a ranking number                                                                      123 select a ranking number                                                                      123
Answer #1

Formula of Payback Period when Same Amount of Cash Inflows in each year:

Initial Investment / Net Annual Cash inflows

In Case of Bono, Payback Period = 11.43 Years (168000/147000)

When Cash Inflows are different in each year then Payback Period Calculation is as follows:

Edge payback Period calculation

Year Cash Inflow/Outflow Cumulative Flow
0 -183750 -183750
1 18900 -164850
2 17850 -145950
3 16800 -127050
4 12600 -108150
5 9450 -89250

=89250/9450 + 5 Years = 9.44 Yeras + 5 Yeras = 14.44 yeras.

Clayton Payback Period calculation

Year Cash Inflow/Outflow Cumulative Flow
0 -210000 -210000
1 28350 -181650
2 24150 -157500
3 22050 -135450
4 13650 -121800
5 12600 -109200

=109200/12600 + 5 Years = 8.66 Years+ 5 years =13.66 years.

Formula For Net Presesnt Value Calculation :Present value of Cash Inflow- PV Of Cash Outflow

Bono's NPV

Year Cash Flows Discounting factor Present Value
0 -168000 1 -168000
1 14700 0.86956 12782.532
2 14700 0.75614 11115.258
3 14700 0.65752 9665.544
4 14700 0.57145 8400.315
5 14700 0.49718 7308.546
49272.195

NPV=49272.195- (168000) =(118727.81)

Edge's NPV

Year Cash Flows Discounting factor Present Value
0 -183750 1 -183750
1 18900 0.86956 16434.684
2 17850 0.75614 13497.099
3 16800 0.65752 11046.336
4 12600 0.57145 7200.27
5 9450 0.49718 4698.351

52876.74

NPV= 52876.74-(183750) =(115123.26)

Claytons NPV

Year Cash Flows Discounting factor Present Value
0 -210000 1 -210000
1 28350 0.86956 24652.026
2 24150 0.75614 18260.781
3 22050 0.65752 14498.316
4 13650 0.57145 7800.2925
5 12600 0.49718 6264.468
71475.8835

NPV=71475.88-(210000) =(96524.12)

Annual Rate of Return formula:Avarage Annual Earnings After Taxes/ Initial Investment*100

Name Bono Edge Clayton
Year Income Income Income
1 14700 18900 28350
2 14700 17850 24150
3 14700 16800 22050
4 14700 12600 13650
5 14700 9450 12600
Total 73500 75600 100800
Avarage 14700 15120 20160
Initial Investent 168000 183750 210000
ARR 8.75 8.23

9.60

As Per Payback Period Bono Project is Recommended.

As per NPV Clayton is Recommended.

As per ARR Clayton Recommended.